# SINDI: An Efficient Index for Sparse Vector Approximate Maximum Inner Product Search

### v2

@@ Abstract.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2026.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/Roxanne0321/vsag/tree/sparse.
1. Introduction
Recently, retrieval-augmented generation (RAG) (Sawarkar et al., 2024; Şakar and Emekci, 2025; Lewis et al., 2020; Su et al., 2025; Quinn et al., 2025; Pogrebinsky et al., 2025) become one of most successful information retrieval framework attracting attention from research communities and industry. Usually, texts are embedded into dense vectors (i.e., no dimension of the vector is zero entry) in RAG, then retrieved through approximate nearest neighbor search (ANNS) on their corresponding dense vectors.
To enhance the RAG framework, researchers find that using sparse vectors retrieval to complement dense vector based RAG can yield better overall accuracy and recall performance. Different from dense vectors, sparse vectors (i.e., only a very small portion of dimensions of spare vectors are non-zero entries) are generated by specific models (e.g., Splade (Formal et al., 2022, 2024; Formal et al., 2021; Lassance and Clinchant, 2022)) to preserve semantic information while enabling precise lexical matching (Ma et al., 2025). In the enhanced RAG framework, dense vectors capture the holistic semantic similarity between texts and sparse vectors ensure exact term recall, therefore resulting in better overall performance. We show the process of the enhanced RAG in the following example:
Figure 1. Example of Dense and Sparse Vector Representations and Inner Product Calculations.
@@ 1. Introduction
Precise lexical matching. In the retriever stage of RAG, queries and documents are compared to select top-$k$ candidates. Dense vectors capture semantic similarity, while sparse vectors support exact term matching. For example, “I love black cats” is tokenized into “i”, “love”, “black”, and “cats”, with “cats” assigned the highest weight (0.8). A query containing “cats” will precisely match documents where this token has a high weight. Challenges in inner-product computation. Dense vectors are stored contiguously, enabling parallel dot-product over consecutive dimensions via SIMD. Sparse vectors typically have very high dimensionality but store only their non-zero entries in a compact format, which leads to two bottlenecks: (1) ID lookup overhead: Matching common non-zero dimensions requires traversing all non-zero entries. Even if only one dimension (e.g., 9999) matches, all entries must be scanned. (2) No SIMD acceleration: Their storage is not dimension-aligned, preventing parallel SIMD processing.
The similarity retrieval problem for sparse vectors is formally known as the Maximum Inner Product Search (MIPS) (Keivani et al., 2018; Pham, 2021; Song et al., 2021; Yan et al., 2018; Zhao et al., 2023), which aims to identify the top-$k$ vectors in a dataset that have the largest inner product value with a given query vector. However, due to the curse of dimensionality (Indyk and Motwani, 1998), performing exact MIPS in high-dimensional spaces is computationally prohibitive. To mitigate this issue, we focus on Approximate Maximum Inner Product Search (AMIPS) (Song et al., 2021; Abuzaid et al., 2019; Neyshabur and Srebro, 2015), which trades a small amount of recall for significantly improved search efficiency.
Many algorithms (Bruch et al., 2024; Zhao et al., 2024; Mallia et al., 2024; PyANNS Contributors, 2025) have been proposed for AMIPS, employing techniques such as inverted index (Bruch et al., 2024), proximity graphs (PyANNS Contributors, 2025), and hash-based (Zhao et al., 2024) partitioning. They improve efficiency by grouping similar vectors into the same partition, thereby reducing the number of candidates examined during query processing.
@@
To address the aforementioned challenges, we propose Sindi, a Sparse Inverted Non-redundant Distance-calculation Index for efficient sparse vector search. The main contributions of this paper are as follows: (i) Value-Storing Inverted Index: Sindi stores both vector identifiers and their corresponding values in the inverted index, enabling direct access during query processing; (ii) Efficient Inner Product Computation: Sindi eliminates redundant overhead in identifying common dimensions and fully exploits SIMD acceleration. By grouping values under the same dimension, Sindi enables batched inner product computation during queries; (iii) Cache-Friendly Design: Sindi reduces random memory accesses by avoiding fetches of original vectors. Instead, it sequentially accesses inverted lists for specific dimensions, thereby lowering cache miss rates. (iv) Vector Mass Pruning: Sindi retains only high-value non-zero entries in vectors, effectively reducing the search space and improving query throughput while preserving accuracy.
Table 1. Comparison to Existing Algorithms.
@@
Distance Complexity $\mathcal{O}\Bigl{(}\frac{\|q\|}{s}\Bigr{)}$ $\mathcal{O}(\|q\|+\|x\|)$ $\mathcal{O}(\|q\|+\|x\|)$
@@
We compare Sindi with several state-of-the-art methods on the MsMarco dataset (8.8M scale) in Table 1. The time complexity of computing the inner product between a query vector $q$ and a document vector $x$ using Sindi is $\mathcal{O}\left(\frac{\lVert q\rVert}{s}\right)$, where the involved symbols are defined in Table 2. In contrast, inverted index and graph-based algorithms have a time complexity of $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$. The detailed derivation is given in § 3.2.
@@
  • We present Sindi, a novel value-storing inverted index described in §3.1 and §3.2, which reduces redundant distance computation and random memory accesses. We further introduce a Window Switch strategy in §3.3 to support large-scale datasets.
  • We propose Vector Mass Pruning in §4 to decrease the search space and improve query speed while maintaining accuracy.
  • We evaluate Sindi on multi-scale, multilingual datasets in §5, demonstrating $4\times\sim 26\times$ higher single-thread QPS than PyANNS and Seismic at over 99% Recall@50, and achieving 8.8M-scale index construction in 60 seconds with minimal cost.
Table 2. Summary of Symbols
@@
$T^{j},T^{j}[t]$ temporary product array on dimension $j$; value at index $t$
@@
2. Preliminaries
2.1. Problem Definition
@@
Let $\mathcal{D}\subseteq\mathbb{R}^{d}$ be a dataset of $d$-dimensional sparse vectors. For any $\vec{x}\in\mathcal{D}$, let $x$ denote its sparse representation, defined as the set of non-zero entries:
$x=\{\,x^{j}\mid x^{j}\neq 0,\;j\in[0,d-1]\,\}.$
Here, $x^{j}$ denotes the value of $\vec{x}$ in dimension $j$. The notation $\lVert x\rVert$ denotes the number of non-zero entries in $x$.
@@
Since the similarity measure in this work is based on the inner product, we formally define its computation on sparse vectors as follows.
@@
Let $\vec{x}_{1},\vec{x}_{2}\in\mathcal{D}$, and let $x_{1}$ and $x_{2}$ denote their sparse representations. Define the set of common non-zero dimensions as
$\Omega(\vec{x}_{1},\vec{x}_{2})=\{\,j\mid x_{1}^{j}\in x_{1}\ \wedge\ x_{2}^{j}\in x_{2}\,\}.$
The inner product between $\vec{x}_{1}$ and $\vec{x}_{2}$ is then given by
$\delta(\vec{x}_{1},\vec{x}_{2})=\sum_{j\in\Omega(\vec{x}_{1},\vec{x}_{2})}x_{1}^{j}\cdot x_{2}^{j}.$
Given the formal definition of the inner product for sparse vectors, we now define the Sparse Maximum Inner Product Search (Sparse-MIPS) task, which aims to find the vector in the dataset that maximizes this similarity measure with the query.
@@
Given a sparse dataset $\mathcal{D}\subseteq\mathbb{R}^{d}$ and a query point $\vec{q}\in\mathbb{R}^{d}$, the Sparse Maximum Inner Product Search (Sparse-MIPS) returns a vector $\vec{x}^{\ast}\in\mathcal{D}$ that has the maximum inner product with $\vec{q}$, i.e.,
(1) $\vec{x}^{\ast}=\underset{\vec{x}\in\mathcal{D}}{\arg\max}\;\delta(\vec{x},\,\vec{q}).$
@@
In practice, $c$-Sparse-MIPS methods can reduce query latency by orders of magnitude compared with exact search, making them preferable for large-scale, real-time applications such as web search, recommender systems, and computational advertising.
For ease of reference, the main notations and their meanings are summarized in Table 2, which will be referred to throughout the rest of the paper.
2.2. Existing Solutions
Representative algorithms for the AMIPS problem on sparse vectors include the inverted-index based Seismic (Bruch et al., 2024), the graph based PyANNS (PyANNS Contributors, 2025). Seismic constructs an inverted list based on vector dimensions. PyANNS creates a proximity graph where similar vectors are connected as neighbors.
Figure 2. The Bottleneck of the Graph Index and Inverted Index During Searching Process.
@@ 2.2. Existing Solutions
Figure 2 illustrates a proximity graph and an inverted index constructed for $\vec{x}_{1}$ to $\vec{x}_{6}$. Consider a query vector $\vec{q}$ with two non-zero entries $q^{1}$ and $q^{2}$. In the proximity graph, when the search reaches $\vec{x}_{4}$, the algorithm computes distances between $\vec{q}$ and all its neighbors, sequentially accessing $x_{1}$, $x_{2}$, $x_{4}$, and $x_{6}$ from memory. In the inverted index, the algorithm traverses the posting lists for dimensions $1$ and $2$, accessing $x_{1}$, $x_{6}$, $x_{2}$, and $x_{4}$. Since vector access during search is essentially random, this incurs substantial random memory access overhead. Moreover, because $\lVert x\rVert$ varies across vectors, the distance computation between $\vec{q}$ and $\vec{x}$ has a time complexity of $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$
Redundant Distance Computations. Sparse vectors incur high distance computation cost due to (i) the explicit lookup needed to identify the common dimensions $\Omega(\vec{x},\vec{q})$ between a document $\vec{x}$ and a query $\vec{q}$, resulting in complexity $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$, and (ii) the inability of existing algorithms to exploit SIMD acceleration for inner product computation. Profiling 6980 queries on the MsMarco dataset (1M vectors) using perf (Linux Kernel Organization, 2025) and VTune (Intel Corporation, 2025) shows that PyANNS spent 83.3% of CPU cycles on distance calculation.
Random Memory Accesses. The inefficiency of memory access in existing algorithms can be attributed to two main factors. First, Seismic organize similar data points into the same partition. To improve accuracy, vectors are replicated across multiple partitions. This replication breaks the alignment between storage layout and query traversal order, preventing cache-friendly sequential access. During retrieval, the index returns candidate vector IDs, which incur random memory accesses to fetch their corresponding data, leading to frequent cache misses. Moreover, $\lVert x\rVert$ varies across sparse vectors, requiring offset table lookups to locate each vector’s data. In our measurements, Seismic averaged 5168 random vector accesses per query (5.1 MB), with an L3 cache miss rate of 67.68%.
3. Full Precision Inverted Index
@@
  • In § 3.1, Sindi constructs a value-based inverted index by storing both vector identifiers and their corresponding dimension values in posting lists. This design eliminates the redundant dimension-matching overhead present in traditional inverted indexes.
  • In § 3.2, Sindi employs a two-phase search process involving product computation and accumulation. By using SIMD instructions in multiplication, it reduces query complexity from $O(\|q\|+\|x\|)$ to $O\Bigl{(}\frac{\|q\|}{s}\Bigr{)}$. This maximizes CPU utilization and improves query throughput. w
  • In § 3.3, to mitigate cache misses caused by random access to the distance array, Sindi introduces Window Switch strategy, which partitions each posting list into fixed-size segments of length $\lambda$ while using a shared distance array. This reduces memory overhead without increasing computation, and both theoretical analysis and experimental evaluation (Figure 5) demonstrate the existence of an optimal $\lambda$ that minimizes memory access overhead.
3.1. Value-storing Inverted Index
@@
Figure 3. Overlap of Inverted List Entries and Common Non-Zero Dimensions in Inner-Product Computation.
@@
Figure 3 illustrates the inverted lists constructed for vectors $x_{1}$ to $x_{5}$. When a query $q$ arrives, it sequentially probes the inverted lists for dimensions 1, 3, and 5. In the base dataset, only the dimensions belonging to $\Omega(\vec{q},\vec{x})$ need to be traversed during the inner product computation. For example, although $x_{4}$ has a value in dimension 2, this dimension is not in $\Omega(\vec{q},x_{4})$ and thus is never accessed during the computation. We further observe that the document identifiers retrieved from the inverted lists overlap exactly with those used to determine the common non-zero dimensions for the inner product. This implies that the products of these non-zero entries can be computed during the document retrieval process itself, thereby ensuring that only dimensions in $\Omega(\vec{q},\vec{x})$ are involved in the inner product computation.
Inspired by these observations, we extend each inverted list $I_{j}$ to store not only the identifier $i$ of the vector $\vec{x}_{i}$, but also the value $x_{i}^{j}$ in dimension $j$. In our notation, we simply write $x_{i}^{j}$ in $I_{j}$ to denote this stored value, with the subscript $i$ implicitly encoding the associated vector ID. This value-storing design eliminates the cost of explicitly locating $\Omega(\vec{x}_{i},\vec{q})$ during inner product computation, as well as the additional random memory access that would otherwise be required to fetch $x_{i}^{j}$ from the original vector.
3.2. Efficient Distance Computation
With the value-storing inverted index, the inner product between a query $\vec{q}$ and candidate vector $\vec{x}$ can be computed without explicitly identifying $\Omega(\vec{q},\vec{x})$. During search, Sindi organizes the computation into two stages: product computation and accumulation.
In the first stage, as each relevant posting list $I_{j}$ is traversed, the products $q^{j}\times x_{i}^{j}$ are computed (using SIMD instructions when possible) and stored sequentially in a temporary array $T^{j}$. Here, $T^{j}$ holds the products for $I_{j}$, with each entry $T^{j}[t]$ corresponding to the $t$-th entry in $I_{j}$. Therefore, $T^{j}$ has the same length as $I_{j}$.
In the second stage, the values in $T^{j}$ are aggregated into a preallocated distance array $A$ of length $\lVert\mathcal{D}\rVert$, where each entry $A[i]$ stores the accumulated score for vector $\vec{x}_{i}$. This arrangement enables $\mathcal{O}(1)$ time per accumulation into $A$, and naturally exploits SIMD parallelism.

### v3

@@ Abstract
I Introduction
Recently, retrieval-augmented generation (RAG) [blendedrag, maxrag, neuripsrag, parametricrag, accrag, enhancerag] has become one of the most successful information retrieval framework attracting attention from research communities and industry. Usually, texts are embedded into dense vectors (i.e., no dimension of the vector is zero entry) in RAG, then retrieved through approximate nearest neighbor search (ANNS) on their corresponding dense vectors [vsag, enhanced_graph, yukun, mingyu].
To enhance the RAG framework, researchers have found that complementing dense vector-based RAG with sparse vector retrieval yields superior accuracy and recall [ma2025lightretriever, sallinen2025mmoremassivemultimodalopen, ahmad2025transformertafsirqias2025, fensore2025evaluatinghybridretrievalaugmented]. Unlike dense vectors, sparse vectors (where only a small fraction of dimensions are non-zero) are generated by specific models (e.g., Splade [distill-splade, splade, efficiencyspalde, towardssplade]) to preserve semantic information while enabling precise lexical matching [ma2025lightretriever]. In the enhanced RAG framework, dense vectors capture holistic semantic similarity, while sparse vectors ensure exact term recall. This synergy translates into significant application-level improvements: in the evaluation of AntGroup production RAG system, integrating sparse vectors into a strong hybrid baseline (BM25 + Dense) boosted Recall@3 by 18.5% (from 55.60% to 74.10%) and Recall@10 by 3.7% (from 85.20% to 88.90%). We illustrate the process of this enhanced RAG workflow in the following example:
Figure 1: Example of Dense and Sparse Vector Representations and Inner Product Calculations.
@@ I Introduction
Precise lexical matching. In the retriever stage of RAG, queries and documents are compared to select top-$k$ candidates. Dense vectors capture semantic similarity, while sparse vectors support exact term matching. For example, “I love black cats” is tokenized into “i”, “love”, “black”, and “cats”, with “cats” receiving the highest weight (0.8). Queries containing “cats” precisely match documents where this token scores high. Challenges in inner-product computation. Dense vectors are stored contiguously, enabling parallel SIMD dot-products. Sparse vectors have high dimensionality but compactly store only non-zero entries, causing two bottlenecks: (1) ID lookup overhead: Matching shared non-zero dimensions requires traversing all entries. Even if only one dimension matches, all must be scanned. (2) No SIMD acceleration: Unaligned storage prevents parallel SIMD processing.
The similarity retrieval problem for sparse vectors is formally known as the Maximum Inner Product Search (MIPS) [ip-1, ip-2, ip-3, ip-4, ip-5], which aims to identify the top-$k$ vectors in a dataset that have the largest inner product value with a given query vector. However, due to the curse of dimensionality [indyk1998approximate], performing exact MIPS in high-dimensional spaces is computationally prohibitive. To mitigate this issue, we focus on Approximate Maximum Inner Product Search (AMIPS) [ip-3, ip-6, ip-7], which trades a small amount of recall for significantly improved search efficiency.
Many algorithms [seismic, sosia, bmp, pyanns] for AMIPS employ techniques like inverted indices [seismic], proximity graphs [pyanns], and hash partitioning [sosia]. They improve efficiency by grouping similar vectors, reducing the candidates examined during queries.
@@
To address the aforementioned challenges, we propose Sindi, a Sparse Inverted Non-redundant Distance-calculation Index for efficient sparse vector search. The main contributions of this paper are: (i) Value-Storing Inverted Index: Sindi stores both vector identifiers and values in the inverted index, enabling direct access during query processing; (ii) Efficient Inner Product Computation: Sindi eliminates redundant overhead in identifying common dimensions and fully exploits SIMD acceleration. By grouping values under the same dimension, Sindi enables batched inner product computation during queries; (iii) Cache-Friendly Design: Sindi reduces random memory accesses by avoiding fetches of original vectors. Instead, it sequentially accesses inverted lists for specific dimensions, thereby lowering cache miss rates. (iv) Vector Mass Pruning: Sindi retains only high-value non-zero entries in vectors, effectively reducing the search space and improving query throughput while preserving accuracy.
TABLE I: Comparison to Existing Algorithms.
@@
Distance Complexity $O\Bigl(\frac{\|q\|}{s}\Bigr)$ $O(\|q\|+\|x\|)$ $O(\|q\|+\|x\|)$
@@
We compare Sindi with several state-of-the-art methods on the MsMarco dataset (8.8M scale) in Table I. Regarding the computational cost, Sindi achieves an amortized per-candidate time complexity of $O\left(\frac{\lVert q\rVert}{s}\right)$ (symbols defined in Table II), establishing a robust performance lower bound across different data distributions. In contrast, traditional inverted index and graph-based algorithms typically scale with $O(\lVert q\rVert+\lVert x\rVert)$. The detailed derivation is given in § III-B.
@@
  • We present Sindi, a novel value-storing inverted index described in §III-A and §III-B, which reduces redundant distance computation and random memory accesses. We further introduce a Window Switch strategy in §III-C to support large-scale datasets.
  • We propose Vector Mass Pruning in §IV to decrease the search space and improve query speed while maintaining accuracy.
  • We evaluate Sindi on multi-scale, multilingual datasets in §V, demonstrating $4\times\sim 26\times$ higher single-thread Queries Per Second (QPS) than PyANNS and Seismic at over 99% Recall@50, and achieving 8.8M-scale index construction in 60 seconds with minimal cost.
TABLE II: Summary of Symbols
@@
$P^{j}$ temporary product array on dimension $j$
@@
II Preliminaries
II-A Problem Definition
@@
Let $\mathcal{D}\subseteq\mathbb{R}^{d}$ be a dataset of $d$-dimensional sparse vectors. For any $\vec{x}\in\mathcal{D}$, let $x$ denote its sparse representation, defined as the set of non-zero entries: $x=\{\,x^{j}\mid x^{j}\neq 0,\;j\in[0,d-1]\,\}.$ Here, $x^{j}$ denotes the value of $\vec{x}$ in dimension $j$. The notation $\lVert x\rVert$ denotes the number of non-zero entries in $x$.
@@
Since the similarity measure in this work is based on the inner product, we formally define it as follows.
@@
Let $\vec{x}_{1},\vec{x}_{2}\in\mathcal{D}$, and let $x_{1}$ and $x_{2}$ denote their sparse representations. Define the set of common non-zero dimensions as $\Omega(\vec{x}_{1},\vec{x}_{2})=\{\,j\mid x_{1}^{j}\in x_{1}\ \wedge\ x_{2}^{j}\in x_{2}\,\}.$ The inner product between $\vec{x}_{1}$ and $\vec{x}_{2}$ is then given by $\delta(\vec{x}_{1},\vec{x}_{2})=\sum_{j\in\Omega(\vec{x}_{1},\vec{x}_{2})}x_{1}^{j}\cdot x_{2}^{j}.$
Given the formal definition of the inner product for sparse vectors, we now define the Sparse Maximum Inner Product Search (Sparse-MIPS) , which finds the vector in the dataset that maximizes this similarity measure with the query.
@@
Given a sparse dataset $\mathcal{D}\subseteq\mathbb{R}^{d}$ and a query point $\vec{q}\in\mathbb{R}^{d}$, the Sparse Maximum Inner Product Search (Sparse-MIPS) returns a vector $\vec{x}^{\ast}\in\mathcal{D}$ that has the maximum inner product with $\vec{q}$, i.e., $\vec{x}^{\ast}=\underset{\vec{x}\in\mathcal{D}}{\arg\max}\;\delta(\vec{x},\,\vec{q}).$
@@
In practice, $c$-Sparse-MIPS methods can reduce query latency by orders of magnitude compared with exact search, making them preferable for large-scale, real-time applications such as web search and recommender systems.
For ease of reference, the main notations and their meanings are summarized in Table II, which will be referred to throughout the rest of the paper.
II-B Existing Solutions
Representative algorithms for the AMIPS problem on sparse vectors include the inverted-index based Seismic [seismic], the graph based PyANNS [pyanns]. Seismic constructs an inverted list based on vector dimensions. PyANNS creates a proximity graph where similar vectors are connected as neighbors.
Figure 2: The Bottleneck of the Graph Index and Inverted Index During Searching Process.
@@ II-B Existing Solutions
Figure 2 illustrates a proximity graph and an inverted index constructed for $\vec{x}_{1}$ to $\vec{x}_{6}$. Consider a query vector $\vec{q}$ with two non-zero entries $q^{1}$ and $q^{2}$. In the proximity graph, when the search reaches $\vec{x}_{4}$, the algorithm computes distances between $\vec{q}$ and all its neighbors, sequentially accessing $x_{1}$, $x_{2}$, $x_{4}$, and $x_{6}$ from memory. In the inverted index, the algorithm traverses the posting lists for dimensions $1$ and $2$, accessing $x_{1}$, $x_{6}$, $x_{2}$, and $x_{4}$. Since vector access during search is essentially random, this incurs substantial random memory access overhead. Moreover, because $\lVert x\rVert$ varies across vectors, the distance computation between $\vec{q}$ and $\vec{x}$ has a time complexity of $O(\lVert q\rVert+\lVert x\rVert)$
Redundant Distance Computations. Sparse vectors incur high distance computation cost due to (i) the explicit lookup needed to identify the common dimensions $\Omega(\vec{x},\vec{q})$ between a document $\vec{x}$ and a query $\vec{q}$, resulting in complexity $O(\lVert q\rVert+\lVert x\rVert)$, and (ii) the inability of existing algorithms to exploit SIMD acceleration for inner product computation. Profiling 6980 queries on the MsMarco dataset (1M vectors) using perf [perf] and VTune [vtune] shows that PyANNS spent 83.3% of CPU cycles on distance calculation.
Random Memory Accesses. The inefficiency of memory access in existing algorithms can be attributed to two main factors. First, Seismic organize similar data into the same partition. To improve accuracy, vectors are replicated across multiple partitions. This replication breaks the alignment between storage layout and query traversal order, preventing cache-friendly sequential access. During retrieval, the index returns candidate vector IDs, which incur random memory accesses to fetch data, leading to frequent cache misses. Moreover, $\lVert x\rVert$ varies across sparse vectors, requiring offset table lookups to locate each vector’s data. In our measurements, Seismic averaged 5168 random vector accesses per query (5.1 MB), with an L3 cache miss rate of 67.68%.
III Full Precision Inverted Index
@@
  • In § III-A, Sindi constructs a value-based inverted index by storing both vector identifiers and their corresponding dimension values in posting lists. This design eliminates the redundant dimension-matching overhead present in traditional inverted indexes.
  • In § III-B, Sindi employs a two-phase search process involving product computation and accumulation. By using SIMD instructions in multiplication, it reduces query complexity from $O(\|q\|+\|x\|)$ to $O\Bigl(\frac{\|q\|}{s}\Bigr)$. This yielding superior CPU utilization compared to state-of-the-art methods and improves query throughput.
  • In § III-C, to reduce cache misses, Sindi uses a Window Switch strategy, partitioning posting lists into fixed-size segments ($\lambda$) sharing a distance array. This minimizes memory overhead, and as shown theoretically and in Figure 5, an optimal $\lambda$ reduces access costs.
III-A Value-storing Inverted Index
@@
Figure 3: Overlap of Inverted List Entries and Common Non-Zero Dimensions in Inner-Product Computation.
@@
Figure 3 illustrates the inverted lists constructed for vectors $x_{1}$ to $x_{5}$. When a query $q$ arrives, it sequentially probes the inverted lists for dimensions 1, 3, and 5. In the base dataset, only dimensions in $\Omega(\vec{q},\vec{x})$ need to be traversed during the inner product computation. For example, although $x_{4}$ has a value in dimension 2, this dimension is not in $\Omega(\vec{q},x_{4})$ and thus never accessed. We further observe that the document identifiers retrieved from the inverted lists overlap exactly with those used to determine the common non-zero dimensions for the inner product. This implies that the products of these non-zero entries can be computed during document retrieval, thereby ensuring that only dimensions in $\Omega(\vec{q},\vec{x})$ are involved in the inner product computation.
Inspired by these observations, we extend each inverted list $I_{j}$ to store not only the identifier $i$ of the vector $\vec{x}_{i}$, but also the value $x_{i}^{j}$ in dimension $j$. In our notation, we simply write $x_{i}^{j}$ in $I_{j}$ to denote this stored value, with the subscript $i$ implicitly encoding the associated vector ID. This value-storing design eliminates the cost of explicitly locating $\Omega(\vec{x}_{i},\vec{q})$ during inner product computation, as well as the additional random memory access needed to fetch $x_{i}^{j}$ from the original vector.
III-B Efficient Distance Computation
With the value-storing inverted index, the inner product between a query $\vec{q}$ and candidate vector $\vec{x}$ can be computed without explicitly identifying $\Omega(\vec{q},\vec{x})$. During search, Sindi organizes the computation into two stages: product computation and accumulation. In the first stage, as each relevant posting list $I_{j}$ is traversed, the products $q^{j}\times x_{i}^{j}$ are computed (using SIMD instructions when possible) and stored sequentially in a temporary array $P^{j}$. Here, $P^{j}$ holds the products for $I_{j}$, with each entry $P^{j}[t]$ corresponding to the $t$-th entry in $I_{j}$. Therefore, $P^{j}$ has the same length as $I_{j}$. In the second stage, the values in $P^{j}$ are aggregated into a preallocated distance array $A$ of length $\lVert\mathcal{D}\rVert$, where each entry $A[i]$ stores the accumulated score for vector $\vec{x}_{i}$. This arrangement enables $O(1)$ time per accumulation into $A$, and naturally exploits SIMD parallelism.